### v1

@@ 6 Conclusion
An extensive review of the use of artificial intelligence (AI) approaches in renewable energy (RE) systems is given in our survey article. Nine AI-based techniques for modern power systems—which address topics including renewable power production, energy forecasting, and system optimization—are found by means of a comprehensive literature study. The reviewed research demonstrate how several AI methods, like as neural networks, decision trees, random forests, and deep learning, are successful in tackling a range of RE management concerns.
The surveyed research demonstrates the potential of AI in enhancing controllability, data handling, cyberattack prevention, and smart grid implementation within the energy industry. Furthermore, the superiority of AI models over conventional approaches is highlighted, particularly in terms of accuracy, efficiency, and adaptability to complex scenarios. Future directions for the integration of AI in renewable energy systems are outlined, emphasizing the need for continued research and development in this rapidly evolving field. Key areas for future exploration include the refinement of AI algorithms, the integration of AI with other emerging technologies such as blockchain and IoT, and the establishment of regulatory frameworks to support AI adoption in the energy sector.
Overall, this survey paper contributes valuable insights into the state-of-the-art AI applications in renewable energy and provides a roadmap for future research endeavors aimed at leveraging AI for sustainable energy transition and environmental conservation.
@@
A fraction of the financial support for this research is provided by DOEd Grant P116Z220008 (1) and NSF Grant No. 2306109. The author(s) presents all viewpoints, discoveries, and conclusions; the sponsor(s) may not necessarily endorse them.
@@
This review paper does not include any original research that involves human beings, animal subjects, or sensitive data.
@@
The papers and data included in this review were obtained from publically accessible and previously published literature.
@@
All authors mentioned in this review study have examined and endorsed the final work for submission to the Machine Learning journal. Every author has made substantial contributions to the research, writing, and revision of the manuscript. All writers have thoroughly reviewed and approved the content included in this review, and they have given their explicit agreement for its publication.
@@
The data and materials utilized are available in the references of this study.
@@
The review is conducted by thoroughly analyzing publically accessible literature, and any mentions of particular approaches or algorithms are properly credited to their original sources. We have made diligent efforts to include explicit citations and references to facilitate readers in locating the source works for additional scrutiny.
@@
Conceptualization: [Parag Biswas], [Angona Biswas]; Methodology: [Abdur Rashid], [Angona Biswas]; Formal analysis and investigation: [Angona Biswas], [Abdullah Al Nasim]; Writing - original draft preparation: [Angona Biswas][Abdullah Al Nasim], [Abdur Rashid],[Parag Biswas]; Writing - review and editing: [Abdur Rashid]; Funding acquisition: [Kishor Datta Gupta]; Resources: [Abdur Rashid], [Parag Biswas]; Supervision: [Kishor Datta Gupta]

### v2

@@ 6 Conclusion
Our study paper provides a thorough analysis of the application of artificial intelligence (AI) methods utilized renewable energy (in short RE) systems. Nine AI-based techniques for modern power systems—which address topics including renewable power production, energy forecasting, and system optimization—are found by means of a comprehensive literature study. The reviewed studies show how various AI techniques, like as deep learning, random forests, decision trees, and neural networks, are effective in addressing a variety of RE management issues.
The surveyed research demonstrates the potential of AI in enhancing controllability, data handling, cyberattack prevention, and smart grid implementation within the energy industry. Furthermore, the superiority of AI models over conventional approaches is highlighted, particularly in terms of accuracy, efficiency, and adaptability to complex scenarios. The paper describes the potential for integrating AI into energy through renewable sources systems in the future and emphasizes the need for further research and development in this rapidly evolving field. Subsequent investigations ought to concentrate on enhancing AI algorithms, merging AI with other state-of-the-art technologies such as blockchain and IoT, and devising legislative structures to promote the application of AI in the field of energy.
Overall, this survey paper contributes valuable insights into the futuristic AI applications in renewable energy as well as provides roadmap for future research endeavors aimed at leveraging AI for sustainable energy transition and environmental conservation.
@@
A portion of the funding for this study comes from NSF Grant No. 2306109 and DOEd Grant P116Z220008 (1). All opinions, findings, and conclusions are presented by the author(s); the the sponsoring organization may not always agree with them.
@@
There isn’t any original study involving humans, animals, or sensitive data in this review publication.
@@
The reviews and data used in this analysis were from previously published, accessible to the public literature.
@@
The final piece of work has been reviewed and approved for obedience to ML journal by every author included in this review research. Each author has contributed significantly to the manuscript’s composition, research, and modification. Each author has carefully examined and approved the material in this analysis, and they have all expressly consented to its publication.
@@
The study’s references provide the information and resources that were used.
@@
The review process involves a comprehensive analysis of publicly available literature, with due acknowledgment given to the original authors of any techniques or algorithms included. We have endeavored to provide clear references and citations so as to aid readers in identifying the original works for closer examination.
@@
Conceptualization: P.B., A.B., M.A.A.N.; Methodology: A.R., A.B., M.A.A.N; Formal analysis and investigation: A.B., M.A.A.N; Writing - original draft preparation: A.B., M.A.A.N., A.R., P.B.; Writing - review and editing: M.A.A.N., A.R.; Funding acquisition: K.D.G., R.G.; Resources: A.R., P.B.; Supervision: K.D.G., R.G. All authors have read and agreed to the published version of the manuscript.